A pacer pencil manufacturer determines that 3 out of every 10000 pacers produced are defective. If 80 000 pencils are produced what is the probability that 20 or more will be defective? Use the continuity correction of adding/subtracting 0.5.
Answer by stanbon(75887)

A pacer pencil manufacturer determines that 3 out of every 10000 pacers produced are defective. If 80 000 pencils are produced what is the probability that 20 or more will be defective? Use the continuity correction of adding/subtracting 0.5.
--------------
mean = np = 80,000(3/10000) = 24
---
std = sqrt(npq) = sqrt(24*(1-0.0003)) = 4.898
------------------------
z(19.5) = (19.5-24)/4.898 = -0.9187
z(80,000.5) = (80,000.5-24)/4.898 = 16328..
P(20<= x <=80,000) = P(-0.9187<= z <= 16328) = 0.8209
===========================================================
Cheers,
Stan H.